OpenStudy (anonymous):

A sound wave with a frequency of 300 hertz is traveling through a medium at a speed of 320 meters/second. What is its wavelength?

OpenStudy (anonymous):

is the ans. 16/5 i.e. 3.2m?

OpenStudy (anonymous):

A sound wave has a wavelength of 0.450 meters. If its speed in cold air is 330 meters/second, what is the wave's frequency? a.0.890hertz b.100hertz c.230hertz d.733

OpenStudy (anonymous):

\[\nu = \frac{ c }{ \lambda }\] \[\lambda = \frac{ c }{\nu }\]

OpenStudy (anonymous):

\[= \frac{ 320 }{ 300 } = 1.06 m\]